OpenStudy (anonymous):

Samantha donated 7 T-shirts to charity and has 14 left. Which equation could be used to find x, the number of T-shirts Samantha had originally? x + 7 = 14 x + 14 = -7 x - 7 = -14 x - 14 = 7

OpenStudy (anonymous):

@flooder

OpenStudy (anonymous):

@BloomLocke367

OpenStudy (anonymous):

@OpenStudyRocks5*

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

@silverr

OpenStudy (bloomlocke367):

is there any that you know you can eliminate?

OpenStudy (anonymous):

c

OpenStudy (bloomlocke367):

okay... any others?

OpenStudy (anonymous):

d

OpenStudy (anonymous):

yes, that once can be eliminated because it ended as a negative. and so can d. so, he gave away 7 from an unknown number. x - 7 there are 14 left afterword = 14 x - 7 = 17

OpenStudy (anonymous):

thats not on there x-7= 17

OpenStudy (anonymous):

its probaly d idk

OpenStudy (anonymous):

or a

OpenStudy (bloomlocke367):

he meant x-7=14... but that's not an option... so you can use that to find the one that will give you the same answer.. make sense?

OpenStudy (anonymous):

yeah so the answer is x-7 = -14

OpenStudy (anonymous):

mabey

OpenStudy (bloomlocke367):

no..... solve for x in the equation that we gave you

OpenStudy (anonymous):

idk sorry

OpenStudy (anonymous):

@lordhelix8th

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

become fan if hlp

OpenStudy (anonymous):

@triciaal

OpenStudy (mathmath333):

\(\large \color{black}{\normalsize \text{let the total number of original T shirst }\hspace{.33em}\\~\\ \normalsize \text{that Samantha had earlier be }\quad x\hspace{.33em}\\~\\ \normalsize \text{she donated 7 Tshirst .}\hspace{.33em}\\~\\ \normalsize \text{So now the total number of T-shirst left will be }\quad (x-7)\hspace{.33em}\\~\\ \normalsize \text{It is given that only 14 T shirst are left , so our equation will be }\hspace{.33em}\\~\\ x-7=14 }\)

OpenStudy (anonymous):

thanks

OpenStudy (mathmath333):

so which option is correct hee @LGumboymath234